OpenStudy (anonymous):

Where does the graph of y = 2x^2 − x − 3 cross the x-axis?

OpenStudy (anonymous):

If the graph crosses the x-axis, then y = 0. 0 = 2x² - 5x + 3 By Factor Theorem: 0 = (2x - 3)(x - 1) Finally, by zero-product identity: 2x - 3 = 0 and x - 1 = 0 x = {3/2,1} I hope this helps!

OpenStudy (anonymous):

yes that is the answer! good luck!